OpenStudy (anonymous):

Convert the equation 12x – 3y = 6 to slope-intercept form.

OpenStudy (anonymous):

slope intercept form is y=mx+c where m is the slope and c the constant

OpenStudy (anonymous):

Okay, I need to know how to convert this equation though. Not just an explanation of what the slope intercept form is.

OpenStudy (amistre64):

slope intercept form is simply: solve for "y"

OpenStudy (amistre64):

how do we get y all alone?

OpenStudy (anonymous):

12x - 3y = 6 -3y = 6-12x y= 6/-3 -12/-3x y=-2 +4x